OpenStudy (simplymarie_x):

Describe how the formulas for a cone, cylinder and pyramid are derived. What are the similarities between the volume of a cone and pyramid?

OpenStudy (michele_laino):

for example, the volume V of a cone whose radius is \(r\), and height is \(h\), is: \[V = \pi {r^2}h\]

OpenStudy (michele_laino):

whereas the volume of a cone, with the same radius and height, is: \[V = \frac{1}{3}\pi {r^2}h\]

OpenStudy (michele_laino):

and we can see that the ratio between such volume is \(1/3\), namely: \[\frac{{{V_{CONE}}}}{{{V_{CYLINDER}}}} = \frac{1}{3}\] provided that both cone and cylinder have the same radius and height

OpenStudy (simplymarie_x):

Wait what? You said volume of a cone twice.

OpenStudy (michele_laino):

I'm saying that if we consider a cone and a cylinder with the same radius of base and the same height, then the volume of cylinder is three time the volume of the cone

OpenStudy (simplymarie_x):

What I mean is you gave me the volume of a cone then said whereas the volume of a cone then you gave me a different formula. Did you make a typo?

OpenStudy (michele_laino):

no, no, I gave the right formulas

OpenStudy (simplymarie_x):

So the volume of a cone has two formulas?

OpenStudy (michele_laino):

the volume V of a cone, is: \[{V_{CONE}} = \frac{1}{3}\pi {r^2}h\]

OpenStudy (michele_laino):

whereas the volume of a cylinder, is: \[{V_{CYLINDER}} = \pi {r^2}h\]

OpenStudy (simplymarie_x):

Okay that's where you confused me, thank you. Because you said volume of a cone gave me the formula then said volume of a cone and gave me a diff formula so you made a typo. It's okay, thought I'd clarify bc it confused me ahah

OpenStudy (michele_laino):

sorry you are right, in the first reply please delete the word "cone" and correct it with the word "cylinder"

OpenStudy (simplymarie_x):

Okay, please keep explaining things ^~^ I'm taking notes as you go

OpenStudy (michele_laino):

I redo, here the corrected first reply: for example, the volume V of a cylinder whose radius is r , and height is h , is: \[{V_{CYLINDER}} = \pi {r^2}h\]

OpenStudy (michele_laino):

here the same second reply: whereas the volume of a cone, with the same radius and height, is: \[{V_{CONE}} = \frac{1}{3}\pi {r^2}h\]

OpenStudy (michele_laino):

finally, here is the same third reply: and we can see that the ratio between such volume is 1/3 , namely: \[\frac{{{V_{CONE}}}}{{{V_{CYLINDER}}}} = \frac{1}{3}\] provided that both cone and cylinder have the same radius and height

OpenStudy (simplymarie_x):

Okay. So does that explain how they are kinda derived? I'm not too sure what derived even means haha.

OpenStudy (michele_laino):

If we know the formula for volume of a cylinder, then I can derive the formula of the volume of a cone with the same radius and the same height, simply dividing by three the formula of the volume of the cylinder

OpenStudy (michele_laino):

namely: \[{V_{CONE}} = \frac{{{V_{CYLINDER}}}}{3}\]

OpenStudy (simplymarie_x):

Okay (-: How would we derive the formula of a pyramid?

OpenStudy (michele_laino):

In a similar way, if I consider a prism with an hexagonal base, whose side is \(l\), and height is \(h\), then its volume \(V\), is: \[{V_{PRISM}} = base\;area \times height\]

OpenStudy (michele_laino):

next, if we consider a pyramid with hexagonal base with the same side \(l\) and the same height \(h\), then its volume, is: \[{V_{PYRAMID}} = \frac{{{V_{PRISM}}}}{3}\]

OpenStudy (michele_laino):

wherein, I repeat, both prism and pyramid have the same height \(h\) and the same side of base \(l\)

OpenStudy (simplymarie_x):

Okay, and last question, what are the similarities between the volume of a cone and pyramid? (-:

OpenStudy (michele_laino):

I think that both formulas have the number \(3\) at the denominator, and the height \(h\) compares, in both formulas, with exponent \(1\)

OpenStudy (simplymarie_x):

Thank you so so much, I appreciate your help!! (-:

OpenStudy (michele_laino):

:)

OpenStudy (michele_laino):

sorry, I meant this: I think that both formulas have the number \(3\) at the denominator, and the height h \(appears\), in both formulas, with exponent \(1\)